OpenStudy (anonymous):

The cost in dollars of a field trip is 64 + 14n where n is the number of people attending what is the cost for 48 people?

OpenStudy (anonymous):

is it 126

OpenStudy (swolesammy):

it is 736

OpenStudy (swolesammy):

14*48=672+64

OpenStudy (anonymous):

64 + (14 * 48) =

OpenStudy (anonymous):

oh ok I got you

OpenStudy (anonymous):

yes @swolesammy is correct.